### v1

# Machine learning analysis of anomalous diffusion

Abstract
The rapid advancements in machine learning have made its application to anomalous diffusion analysis both essential and inevitable. This review systematically introduces the integration of machine learning techniques for enhanced analysis of anomalous diffusion, focusing on two pivotal aspects: single trajectory characterization via machine learning and representation learning of anomalous diffusion. We extensively compare various machine learning methods, including both classical machine learning and deep learning, used for the inference of diffusion parameters and trajectory segmentation. Additionally, platforms such as the Anomalous Diffusion Challenge that serve as benchmarks for evaluating these methods are highlighted. On the other hand, we outline three primary strategies for representing anomalous diffusion: the combination of predefined features, the feature vector from the penultimate layer of neural network, and the latent representation from the autoencoder, analyzing their applicability across various scenarios. This investigation paves the way for future research, offering valuable perspectives that can further enrich the study of anomalous diffusion and advance the application of artificial intelligence in statistical physics and biophysics.
keywords:
anomalous diffusion, machine learning, single trajectory characterization, representation learning
1 Introduction
Anomalous diffusion[1, 2, 3], which refers to any type of diffusion that deviates from classical Brownian motion, is a fundamental phenomenon across a wide range of scientific disciplines, such as physics[4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], chemistry[15, 16, 17], biology[18, 19, 20, 21, 22, 23, 24], and finance[25, 26, 27, 28]. As shown in figure 1(a), numerous experiments and observations have demonstrated that the occurrence of anomalous diffusion is ubiquitous across various scales and systems[29]. For instance, nanoparticles navigating through polymer networks with semiflexible strands perform enhanced heterogeneous diffusion [figure 1(b)][13, 14], while intracellular transport dynamics often exhibits distinct characteristic of anomalous diffusion [figure 1(c)][24]. Even in ecological contexts, the seasonal migrations of species like springboks demonstrate non-standard movement patterns on a large scale [figure 1(d)][28]. To comprehensively understand the mechanisms underlying these phenomena, the characterization and analysis of anomalous diffusion are crucial and increasingly demanded, providing key insights in the research of complex systems[30, 31, 32, 33, 34].
The analysis of anomalous diffusion primarily emphasizes its distinctions from Brownian diffusion, with several characteristics typically identified. One notable characteristic is the nonlinearity of mean squared displacement (MSD):
${\rm MSD}(t)\sim t^{\alpha},$ (1)
where the diffusion exponent $\alpha$ categorizes the diffusion as either superdiffusive ($\alpha>1$) or subdiffusive ($\alpha<1$). Another distinct aspect involves the stochastic process in anomalous diffusion, which contrast significantly with the Wiener process associated with Brownian motion. For this reason, anomalous diffusion trajectories often exhibit long-range correlations and memory effects, or demonstrates the “anomalous yet Brownian" diffusion characteristic [35, 36, 37, 38, 39, 40]. As a response, numerous theoretical models used to describe the evolution of anomalous diffusion have been proposed, such as the continuous-time random walk (CTRW)[41, 42] and fractional Brownian motion (FBM)[43]. CTRW stands out for its integration of random steps with variable waiting times, effectively modeling memory and non-Markovian effects. FBM extends traditional Brownian motion by incorporating a Hurst exponent, allowing it to describe random processes exhibiting persistent or anti-persistent characteristics. A third key distinction between anomalous diffusion and Brownian motion arises in complex environments, where a random walker may exhibit varying diffusion states over time due to intrinsic or extrinsic changes, leading to heterogeneous diffusion dynamics[44, 45, 46, 13, 14, 47, 48, 49, 50, 51, 52, 53, 54, 55]. For instance, the Hfq protein in Escherichia coli cells can continuously transition among three diffusion states with different diffusion coefficients and mean dwell times [45]. Similarly, in the microtubule network of DU145 cells [44], the transports of endosomes can vary among Lévy walk (LW) [56] and Brownian motion. These complexities emphasize the need for advanced methods to decode the diffusion data and unravel the diverse nature of anomalous diffusion.
Figure 1: (a) Illustrative images demonstrate anomalous diffusion across various length scales (top panel), and different dimensions (bottom panel), from [29]; (b) Enhanced heterogeneous diffusion can be identified for nanoparticles in a semiflexible polymer network, from [13, 14]; (c) Intracellular transport dynamics in early apoptotic cells, where trajectories of endocytic vesicles display typical anomalous diffusion properties, from [24]; (d) GPS-tracking movement data of springboks in Namibia during the wet season (top) and the dry season (bottom), both showing non-standard movement patterns on a large scale, from [28].
To address the challenges of analyzing anomalous diffusion, statistical methods[1, 57, 58, 30, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72] are extensively used before the advent of machine learning techniques. These conventional approaches primarily relies on the construction of parametric features that effectively characterize the statistical properties of diffusion trajectories. For example, one key technique is the calculation of MSD, which can be used to infer the diffusion coefficients and diffusion exponents. Another is the self-part of the van Hove function[69] that quantifies the displacement distribution of a random walker. The non-Gaussianity of this function allows for a deeper understanding of the deviations from classical Brownian motion in anomalous diffusion behaviors. The strong interpretability of these statistical methods has led to their widespread use in the diffusion research, as they provide clear insights into the underlying mechanisms of diffusion processes.
However, inherent limitations of traditional statistical methods hinder their application, especially in modern research contexts. For example, these methods typically require long trajectories or large datasets to yield statistically reliable results. This requirement makes them less effective when dealing with short tracks or small datasets, which are often encountered in experimental settings. Another issue is their frequently inadequate accuracy, particularly in distinguishing between closely related diffusion behaviors under complex conditions. This is partly due to the reliance on assumptions that may not hold true across different systems or scales, such as the presumption of ergodicity or the application of homogeneous diffusion models to inherently heterogeneous environments. Moreover, these statistical methods are notably vulnerable to noise. This susceptibility can lead to potential misinterpretations of diffusion behaviors in trajectories with low signal-to-noise ratios, especially when experimental conditions are not ideal[73]. These limitations underscore the urgent demand for developing new analytical methods that can more effectively handle the diversity and complexity of anomalous diffusion data.
On the other hand, as an essential branch of artificial intelligence, machine learning techniques have undergone a rapid development in recent years[74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88]. On the basis of big data, ML models excel in directly identifying patterns, making predictions, and informing decisions. In particular, ML is tremendously powerful in processing time-series data [85, 86, 87, 88], which is critical for fields ranging from financial forecasting[85] to health monitoring[87] and environmental sensing[88]. The significance of ML in time-series analysis lies in its ability to model and predict complex sequential information without explicit programmatic definitions. ML algorithms can learn from historical data, allowing them to capture temporal dynamics and dependencies that are often missed by traditional statistical methods. This capability enables them to perform tasks such as anomaly detection, trend forecasting, and event prediction with high accuracy. Given the inherent nature of anomalous diffusion trajectories, which can be regarded as complex time-series data, ML is well-positioned to provide substantial analytical insights into these phenomena. It is foreseeable that ML algorithms will adeptly decode the subtle nuances in diffusion behavior by learning from the temporal sequence of data points, capturing long-range dependencies and patterns that may not be immediately obvious. Consequently, by leveraging ML’s robust capabilities in time-series analysis, researchers can expect to obtain advanced tools for the analysis of diffusion processes.
Building upon these expectations, this review systematically summarizes the integration of ML techniques in the analysis of anomalous diffusion in recent years. The paper is organized as follows: We start with an examination of single trajectory characterization via ML in section 2, which focuses on the inference of diffusion parameters and segmentation of heterogeneous diffusion dynamics. ML competitions regarding anomalous diffusion analysis are also introduced in section 2. These competitions have catalyzed the development of a variety of high-performance ML algorithms specifically tailored for analyzing anomalous diffusion. Section 3 provides a comprehensive overview of research related to the representation learning of anomalous diffusion, discussing how to effectively represent the dynamics of anomalous diffusion. Finally, we summarize our findings and suggest promising research avenues in this domain in section 4.
2 Single Trajectory Characterization via Machine Learning
The characterization of single trajectory properties is a fundamental step in providing detailed insights into the underlying physics of anomalous diffusion. Currently, the primary focus of single trajectory characterization centers around two critical tasks: the inference of diffusion parameters and the segmentation of heterogeneous diffusion dynamics, also known as trajectory segmentation. In this section, we will detail the applications of machine learning methods to these two tasks, along with the Anomalous Diffusion Challenge [89, 29], which is a competition designed to benchmark the capabilities of various machine learning models against these tasks.
2.1 Inference of diffusion parameters
Accurately obtaining key diffusion parameters of trajectories is paramount in single trajectory characterization. Among the myriad diffusion parameters, the most representative parameters include the diffusion coefficient ($D$), the diffusion exponent ($\alpha$), and the theoretical diffusion model. The coefficient $D$ is a parameter that quantifies the diffusion speed, which can reflect properties such as the viscosity and temperature of the medium. The exponent $\alpha$ appears in the relationship between the MSD and time, revealing whether the diffusion is normal, subdiffusive, or superdiffusive. The diffusion model, such as CTRW and FBM, refers to the mathematical and physical framework used to theoretically describe the diffusion process. Based on the combined analysis of these parameters, researchers can understand the motion behaviors and dynamic mechanisms underlying the trajectory data.
Therefore, before the utilization of machine learning, numerous methods have been developed to infer these diffusion parameters. A classic example is the use of time-averaged mean squared displacement (TA-MSD) to determine the diffusion exponent $\alpha$ based on Eq. (1). Another instance is the analysis of power spectral density (PSD), which provides the frequency-dependent information of the trajectories [57, 58, 30, 59]. In addition, velocity autocorrelation function (VACF) [60], fractional integral moving average (FIMA) process [61, 62, 63] and $p$-variation test[64] are also extensively employed when calculating the diffusion parameters. These traditional methods offer straightforward and highly interpretable analytical approaches that help us uncover the physical information embedded in diffusion trajectories.
However, intrinsic limitations are inevitable for statistics-based techniques. These traditional methods are generally sensitive to outliers and anomalies, and typically have a narrow scope when performing parameter inference. Each method is usually capable of inferring only one specific diffusion parameter, or distinguishing only two or few diffusion models. More importantly, with the rapid development of single particle tracking (SPT) technology [90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108], a growing array of diverse diffusion datasets has become available. Traditional statistics-based methods often struggle to effectively manage this complexity. These limitations underscore the urgent need to develop methods with robust generalization capabilities and the ability to handle complex datasets effectively.
To address this challenge, machine learning techniques have been introduced to the inference of diffusion parameters due to its powerful ability to handle complex data in recent years. In particular, inferring diffusion parameters essentially constitutes a regression or classification problem, fitting well within the standard supervised learning paradigm. For clarity in our discussion, we have organized these ML approaches into two primary categories in this section: classical machine learning and deep learning. Within classical machine learning, we highlight significant contributions using techniques such as feature-based models[109, 110, 111, 112] and Bayesian methods[113, 114, 115, 116, 117, 118]. Regarding deep learning, we focus on innovative applications that employ deep neural networks, including convolutional neural networks [119, 120, 121, 122, 123, 124, 125, 126, 127], recurrent neural networks[128, 129, 130, 131, 132], graph neural networks[133, 134, 135], to analyze diffusion trajectories and infer the parameters.
2.1.1 Classical machine learning
Classical machine learning encompasses a range of algorithms and techniques that predate the extensive use of deep neural networks. These methods are grounded in established statistical learning theories and are characterized by their simplicity, interpretability, and lower computational requirements. For the analysis of anomalous diffusion, classical machine learning primarily relies on feature-based models or Bayesian approaches to determine diffusion parameters.
As highlighted in figure 2(a), feature-based models operate on manually extracted features from diffusion trajectories and utilize classic machine learning algorithms such as decision trees and multilayer perceptrons for parameter inference. By harnessing handcrafted features that capture the essential characteristics of the diffusion process, such as fractal dimension, maximal excursion, and mean gaussianity [70, 71, 72], these models provide a structured way to analyze and interpret diffusion data. Typical related works include the following:
Figure 2: (a) Schematic flowchart of the feature-based methods: a set of features is extracted from raw trajectories and used as the input for feature-based machine learning models, from [73]; (b) A workflow of random forest algorithm for the single trajectory characterization tasks, from[109]; (c) Schematic diagram of the Markovian approximation of discrete-time Lévy walks (top); comparisons between Bayesian inference and MSD fitting across different signal-to-noise ratio conditions when predicting diffusion exponents (bottom), from [115].
  • A study by G. Muñoz-Gil et al. [109] employs the random forest (RF) algorithm for the single trajectory characterization tasks, which is a classical machine learning method by aggregating results of multiple decision trees through voting or averaging As shown in figure 2(b), upon the supervision of labeled simulated trajectories from various diffusion models, the RF model can be trained to infer diffusion exponent and theoretical model after preprocessing and feature extraction of these trajectories. The trained model is then applied to experimental data, including the movement of mRNA molecules in bacterial cells and membrane receptor diffusion in living cells. The results show that the RF model achieves performance superior to traditional methods, particularly in the analysis of nonergodic diffusion processes.
  • The work by P. Kowalek et al. [110] utilizes the gradient boosting algorithm to analyze the diffusion trajectories. This algorithm is an ensemble machine learning technique that progressively refines models by addressing errors from previous iterations. Each subsequent model is specifically trained to minimize the residuals left by its predecessors, culminating in a prediction that represents a weighted amalgamation of all successive models. In particular, the authors choose XGBoost [79] in this work, which is a decision-tree-based version of gradient boosting algorithm. Initially, the study involves extraction and selection of diverse features from raw diffusion trajectories, utilizing the XGBoost to categorize different diffusion behaviors. This process is iteratively refined by integrating new features, significantly enhancing the classifier’s accuracy. The robustness of this method is further validated by its application to complex diffusion patterns, illustrating its potential in providing insightful analysis of anomalous diffusion dynamics from various perspectives.
  • Another application of classical machine learning is the use of extreme learning machine (ELM) in C. Manzo’s work [111]. This method is tailored for single-layer feedforward networks, simplifying the learning process by randomly assigning weights between the input and hidden layers. In this study, ELM is integrated with feature engineering to analyze anomalous diffusion trajectories. The input to the ELM model includes features such as time-normalized mean absolute displacement, mean square displacement, and cumulative square displacement. These features are efficiently processed by ELM to infer critical diffusion parameters like diffusion exponent and diffusion model. The model performance underscores the effectiveness of ELM in handling complex, feature-rich datasets, demonstrating its potential to deliver rapid and accurate analysis of diffusion behaviors.
In practice, the effectiveness of these feature-based models heavily depends on the statistical features extracted from trajectory data, such as TA-MSD, displacement moments, or other engineered features that characterize the diffusion dynamics, which will be discussed in detail in section 3.1. As these features typically possess strong statistical or physical significance, they provide feature-based methods with enhanced interpretability compared to deep learning. However, selecting the optimal combination of features is challenging, which requires deep domain expertise and extensive experimental analysis to identify the most significant characteristics of diffusion properties. In this regard, H. Loch-Olszewska et al. underscores the impact of feature choice on machine learning classification of anomalous diffusion [112]. The study shows that while combining multiple feature sets can enhance the performance of classifier, no single feature set can universally excel across all diffusion models. This variability highlights the profound impact of feature selection, and also suggests that feature-based machine learning methods may have limitations in their universality for the analysis of anomalous diffusion.
Beyond the feature-based methods, Bayesian inference represents another powerful classic machine learning strategy [113, 114, 115, 116, 117, 118], renowned for its robust interpretability and analytical depth in parameter inference for anomalous diffusion. Bayesian methods offer a probabilistic perspective that enables the integration of prior knowledge with observed data, facilitating a dynamic update of our understanding as new data becomes available. This probabilistic framework employs posterior distributions derived from Bayesian updates to quantify the uncertainty in parameter estimates effectively. By updating the prior beliefs with likelihoods calculated from new data, Bayesian inference continually refines the estimates of model parameters. This approach not only aids in selecting the model that best explains the data but also enhances the understanding of the uncertainty and variability inherent in the model predictions.
One representative work of Bayesian inference for anomalous diffusion is conducted by S. Park et al. [115], which enables an efficient parameter estimation of Lévy walks approximated by the hidden Markov model [figure 2(c)]. The Bayesian approach in this work allows for the inference of the power-law exponent of flight-time distribution and the classification of different diffusion models. This method excels in handling noise and incomplete data, providing robust estimations even when traditional methods like MSD fitting fail. Another work by J. Krog et al. [117] combines Bayesian inference with an information content model check to analyze stochastic processes described by Langevin equations. The authors apply this approach for SPT data to estimate model parameters and compare different candidate models. The results provide not only the most probable parameters but also their uncertainties. In addition, H. Seckler et al. have combined the Bayesian framework with deep learning for error estimation in diffusion trajectory analysis [114]. These studies demonstrate that Bayesian inference provides a flexible and powerful paradigm capable of integrating various tools to achieve the analysis of anomalous diffusion.
2.1.2 Deep learning
While classical machine learning techniques have demonstrated substantial utility in analyzing anomalous diffusion, the advent of deep learning offers a transformative approach due to its inherent capability to manage big data with complex patterns. Unlike feature-based methods that require manual feature extraction, deep learning strategies prefer integrating raw trajectory data directly into deep neural network architectures. This shift enables a more granular understanding of diffusion processes by leveraging minimal preprocessing, maintaining focus on the intrinsic temporal sequences of trajectory data. Building on this advanced methodology, deep learning models such as convolutional neural network (CNN) [82, 119, 120, 121, 122, 123, 124, 125, 126, 127], recurrent neural network (RNN) [83, 128, 129, 131, 130, 132], and graph neural network (GNN) [84, 133, 134, 135] are at the forefront of analyzing anomalous diffusion.
CNN stand as a fundamental neural network architecture in deep learning, particularly renowned for its ability to process spatial and temporal data like images and time series. It excels in autonomously extracting hierarchical features from raw data, which is crucial for comprehensive analysis across various scales. Pioneering this approach in anomalous diffusion, the study by N. Granik et al.[120] shows the application of CNNs in the classification and parameter estimation of single particle trajectories. As shown in the top of figure 3(a), the authors transform the 2D trajectory data into a numerical matrix as the input. Subsequently, by employing a CNN where the layers incrementally increase dilation factors, this model adeptly predicts the motion types and captures correlations of trajectory data across extensive temporal spans. Another study conducted by E. A AL-hada et al.[121] explores the adaptability of CNNs in discerning complex stochastic processes, where the input is the time-evolution image of a trajectory [bottom of figure 3(a)]. This research utilizes an array of pre-trained CNN models, including ResNet-18, ResNet-50 [74], and GoogLeNet [75], to delve into various stochastic categories like CTRW, FBM, LW, and multi-state processes. The results demonstrate the superior capability of CNNs in classifying these stochastic behaviors with high accuracy. In addition, the use of 1D CNN [76, 123, 122, 126, 119] enables the model without the need for additional trajectory preprocessing, enhancing the CNN’s ability to handle raw trajectory data directly. These works collectively illustrate the effectiveness of CNNs in inferring parameters of anomalous diffusion.
Figure 3: (a) Schematic diagrams of single trajectory characterizations using CNNs, from [120, 121]; (b) The workflow (top) and model performance (bottom) of RANDI model, which has a dual-layer LSTM structure, from [129]; (c) Workflow (left) and performance analysis (right) of WADNet, which is combined by a WaveNet encoder and a 3-layer LSTM network, from [123].
Compared to the diverse preprocessing requirements of CNNs for trajectory data, RNNs are inherently suited for time-series analysis, allowing them to process raw trajectory data with minimal preprocessing. On the other hand, the memory capabilities of RNNs, especially long short-term memory (LSTM) networks, enable them to effectively analyze the long-range correlations within trajectories. A notable application of RNNs in this domain is the work by S. Bo et al. [128], where the LSTM network is leveraged to infer the diffusion exponent from experimental trajectory data of sub-diffusive colloids and super-diffusive microswimmers. This work underscores the ability of RNNs to process the intricate dynamics of such systems, especially in environments with speckle light fields where traditional methods falter. Further expanding on this, the RANDI model, developed by A. Argun et al. [129], incorporates a dual-layer LSTM structure [figure 3(b)], refining the model’s ability for the inference of diffusion parameters. These studies [128, 129, 131, 130, 132] illustrate RNNs’ potential to provide deep insights into the underlying physics of anomalous diffusion.
Moreover, following the successful applications of CNNs and RNNs in anomalous diffusion, researchers have begun experimenting with hybrid architectures that combine different types of machine learning models to achieve better performance. For instance, as shown in figure 3(c), D. Li et al.[123] propose a deep learning architecture called WADNet, which has an excellent ability for the diffusion parameter inference. This model combines an improved WaveNet [80] encoder (a specialized 1D CNN) with a 3-layer LSTM network to infer diffusion exponents and classify diffusion models. In particular, WADNet achieves outstanding performance in the $1^{\rm st}$ Anomalous Diffusion Challenge [89], surpassing the first places in the competition leaderboard across all tasks and dimensions. In addition to this, there have been many combination attempts, such as combining ResNet with XGBoost [124], CNN with bi-LSTM [125], and Bayesian inference with deep networks [114]. These strategies not only integrate the strengths of various networks but also compensate for the shortcomings of individual models, making them a more effective method for the analysis of anomalous diffusion.
Figure 4: (a) Representative example of graphs associated with a short diffusion trajectory, from [133]; (b) Workflow of the GNN for processing diffusion trajectories, from [133]; (c) Schematic diagram showing the spatiotemporal graph representation of trajectories using MAGIK, from [135]; (d) Comparative analysis of trajectory linking results using MAGIK on HeLa cell video data with ground truth trajectories, from [135].
Furthermore, GNN offers a distinct advantage for analyzing anomalous diffusion by leveraging their unique ability to process graph-structured data. This capability is essential when representing diffusion trajectories as complex graphs, where nodes symbolize discrete trajectory points and edges encapsulate the temporal or spatial relationships between these points. GNNs excel in handling these non-Euclidean structures through a message-passing mechanism that integrates multi-scale information across various nodes. By employing iterative aggregation, GNNs capture both immediate and extended temporal dependencies, providing a deep understanding of the complex relationships inherent in diffusion processes. For instance, in the work by H. Verdier et al.[133] , the GNN model is structured to transform each trajectory into a graph, as illustrated in figures 4(a) and 4(b). The input graph is processed through several graph convolution operations that project it into a latent space [figure 4(b)]. This representation facilitates the detailed estimation of the anomalous exponent and classifies the type of random walk effectively. Another example is the development of MAGIK framework by J. Pineda et al. [135]. MAGIK utilizes a geometric deep learning approach that models motion through graph representations, employing attention-based components for dynamic estimation [figure 4(c)]. This framework showcases flexibility across diverse biological experiments [figure 4(d)], proving GNNs as a robust tool for spatiotemporal analysis without the need for explicit trajectory linking. These robust applications of GNNs [133, 134, 135] highlights their potential to significantly enhance our understanding of complex stochastic processes.
2.1.3 Further discussions
While both classic machine learning and deep learning have demonstrated outstanding performance in the inference of diffusion parameters, there are critical considerations that need to be addressed for these methods to reach their full potential. First, the application of machine learning in this field primarily falls within the paradigm of supervised learning, where the task is either a classification or regression problem. Therefore, large and high-quality datasets with accurate labels are required when training these models. However, obtaining labels from experimental data can be challenging. To address this issue, the main current approach is the simulation-based method [109, 110, 111, 112, 114, 120, 121, 122, 119, 123, 126, 124, 125, 128, 129, 130, 131, 132, 133, 135, 134, 136]. This method uses various known mathematical diffusion models, such as CTRW, FBM, and LW, to simulate trajectories. These trajectories can be labeled based on the diffusion models and corresponding simulation parameters. After training the machine learning models on simulated trajectories, these models can be subsequently applied to the analysis of real experimental data. Although this method has shown excellent performance so far, it faces an inherent challenge due to the natural differences between simulated and experimental data. On one hand, experimental trajectories often exhibit higher levels of noise, stochasticity, and complexity compared to their simulated counterparts. On the other hand, many characteristics of experimental trajectories cannot be fully captured and simulated by known diffusion models. This clearly limits the generalizability of machine learning models trained solely on synthetic data. Thus, the improvement and optimization of the simulation-based method will be a critical issue for future research.
Second, each of classical machine learning and deep learning has its own strengths and limitations when applied to the inference of diffusion parameters. Understanding these differences is crucial for selecting the appropriate method based on the research objectives. In the following, we compare the characteristics of these two types of machine learning models from three perspectives: input form, interpretability, and nonlinear modeling capabilities.
  1. 1.
    The forms of input data significantly diverge between classical machine learning and deep learning techniques. Classical machine learning depends heavily on preprocessing and feature engineering. This task often necessitates considerable domain knowledge and can be labor-intensive. The effectiveness of these models is largely contingent upon the quality and appropriateness of the chosen features, making them particularly effective when the features are tailored to the anomalous diffusion. Conversely, deep learning methods excel in processing raw data, largely circumventing the need for manual feature engineering. These models can autonomously learn to identify relevant features during their training processes. This direct data consumption enables deep learning models to be more adaptable and scalable, advantageous in complex scenarios or when the critical features of the data are unknown.
  2. 2.
    The interpretability of models represents a crucial distinction between classical machine learning and deep learning approaches. Classical techniques like decision trees and random forests offer higher transparency due to their reliance on explicitly defined features. As stated before, these crafted features are selected with substantial domain knowledge. They often root in well-understood physical or statistical properties of the diffusion process, such as TA-MSD and VACF. This level of clarity is invaluable, especially in scientific research where understanding the underlying mechanisms is as crucial as the predictive performance of the model. On the contrary, deep learning models are often criticized for their “black box" nature, as their internal workings remain largely opaque. Nevertheless, it is possible to gain insights into the role of each component within deep neural networks through methods such as ablation studies [137, 138]. These techniques can help elucidate the underlying mechanisms by which neural networks infer diffusion parameters.
  3. 3.
    The capacity to model nonlinear relationships is another essential distinction between classical machine learning and deep learning. Classical approaches often rely on linear assumptions or straightforward nonlinear transformations, thereby emphasizing the importance of feature engineering. By contrast, deep learning is inherently adept at capturing complex nonlinear interactions due to its multilayered architecture. This profound capability to model nonlinearities is crucial when dealing with anomalous diffusion. This is because the nonlinear information within anomalous diffusion trajectories, such as long-range correlations and frequency space information, is often implicit and challenging to fully uncover through manual feature extraction. In particular, deep learning can autonomously discover these hidden complex nonlinear patterns. Although these patterns remain implicit within neural networks, this does not hinder deep learning from being a potent tool in characterizing anomalous diffusion.
2.2 Segmentation of heterogeneous diffusion dynamics
As highlighted in the introduction, a significant distinction between anomalous diffusion and Brownian motion emerges in complex environments, where a random walker may exhibit varying diffusion states over time due to intrinsic or extrinsic changes, leading to heterogeneous diffusion dynamics [44, 45, 46, 13, 14, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Compared to single-state trajectories with fixed properties, heterogeneous dynamics involve transitions among multiple diffusion states. This makes the accurate analysis of multi-state trajectory data much more complex and thus poses significant challenges in the single trajectory characterization. Segmentation of heterogeneous diffusion dynamics is therefore a crucial task in the analysis of anomalous diffusion. This task involves two primary subtasks:
  1. 1.
    Identifying transition points where changes in the diffusion state occur. We call these transition points as changepoints in this work.
  2. 2.
    Determining the diffusion parameters for each segmented sub-trajectory. For convenience, we use the term segments to refer to these segmented sub-trajectories.
The combination of these two complex subtasks leads to a diverse range of techniques for segmenting trajectories of heterogeneous diffusion dynamics. In this article, we primarily introduce two mainstream types of segmentation approaches: the sliding window and point-wise methods.
Figure 5: (a) Illustration of three segmentation steps in the DC-MSS algorithm, from [139]; (b) The combination of sliding window method and BPNN for trajectory segmentation (top) and corresponding performance analysis (bottom) , from [140]; (c) Workflow of Deep-SEES for trajectory segmentation where the sliding window and LSTM-VAE network are utilized, from [141].
2.2.1 Sliding window method
As the name suggests, the sliding window method involves segmenting a trajectory progressively using a sliding window algorithm [142]. The window length is predefined, dividing the trajectory into several segments of equal length. After that, a classifier is employed to predict the diffusion model associated with each segment. This classifier can be either a traditional statistical method or a machine learning model. By detecting the abrupt changes in predictions for consecutive segments, the changepoints of diffusion states can be identified.
Before the application of machine learning to the analysis of anomalous diffusion, researchers have already employed the sliding window method to segment multi-state diffusion trajectories. A notable example is the divide-and-conquer moment scaling spectrum (DC-MSS) approach developed by A. R. Vega et al.[139]. As depicted in figure 5(a), DC-MSS begins with an initial segmentation using an 11-frame sliding window to identify potential changepoints. This is followed by the classification of these segments through moment scaling spectrum analysis, which categorizes them into different motion types. The final step involves refining the segmentation by merging and reclassifying adjacent segments to ensure accurate changepoint identification. This method is tested on tracking data of the transmembrane protein CD44 on the surface of macrophages, revealing transitions between various motion types, including free diffusion, confined diffusion, directed diffusion, and immobilization.
While the sliding window method permits trajectory segmentation using traditional statistical approaches, these approaches often fall short when faced with more complex trajectory data. Consequently, integrating machine learning techniques with the sliding window method has enabled the creation of models that possess enhanced segmentation capabilities. An illustrative instance of this integration is the algorithm developed by P. Dosset et al.[140], which employs a back-propagation neural network (BPNN) in conjunction with a 31-frame sliding window to classify diffusion models. As shown in figure 5(b), BPNN performs localized MSD analysis on segments delineated by the sliding window, enabling the detection of transitions among distinct diffusion models such as Brownian, confined, and directed motions. This algorithm has been tested on experimental trajectories of transmembrane proteins and viral particles, validating its capacity to accurately segment multi-state diffusion trajectories.
Another application of this technique is the Deep-SEES method introduced by Y. Zhang et al.[141]. Utilizing a sliding window approach, this method segments trajectories for subsequent analysis with a long short-term memory with variational autoencoder (LSTM-VAE) network, as illustrated in figure 5(c). This network excels at extracting latent features and classifying dynamic states, addressing challenges like noise and undersampling effectively. This segmentation capability is crucial for identifying rare events, such as when nanoparticles cross the membrane, which may be hidden within complex or noisy trajectories. Specifically, in active enzyme systems, Deep-SEES adeptly segments the heterogeneous dynamics of gold nanorods. Similarly, in liquid-liquid phase separation systems, it accurately captures the distinct rotational behaviors of nanoparticles. Additionally, other machine learning models such as support vector machine (SVM) [143] and K-means clustering [144, 145], when integrated with the sliding window method, have also demonstrated effectiveness in trajectory segmentation. For a detailed overview of such integrations and their features, refer to the selected examples listed in Table 1. These integrations further broaden the application spectrum of sliding window methods, offering diverse analytical tools capable of addressing the segmentation of heterogeneous diffusion dynamics.
2.2.2 Point-wise method
While the sliding window method offers a systematic approach to segmenting trajectories, it inherently suffers from limitations, particularly regarding the predefinition of window length. The choice of window length is often a trade-off: larger windows might mask rapid transitions between states, while smaller ones could increase noise and lead to misclassifications. These constraints become more pronounced when dealing with complex and heterogeneous diffusion dynamics, where transitions between states do not necessarily follow uniform lengths.
Figure 6: (a) Schematic diagram of the STEP model, from [146]; (b) Segmentation results of diffusion coefficients given by STEP for experimental trajectories of integrin $\alpha 5\beta 1$ (left) and the pathogen-recognition receptor DC-SIGN (right), from [146]; (c) Workflow of the U-AnDi model, from [138]; (d) Segmentation performance of U-AnDi on the diffusion of transmembrane proteins on cell membrane surfaces, from [138].
Table 1: Overview of selected methods for the segmentation of anomalous diffusion.
Method Name Method Type Learning Type Sim.-Based Seg. Target
Divide-and-conquer moment scaling spectrum (DC-MSS)[139] Sliding window Stat. No model
Sliding-window time-averaged MSD [144] Sliding window Stat. No $\alpha$
Recurrence analysis with sliding window segmentation[145] Sliding window Stat. No model
Probability density function and autocorrelation function analysis[147] Sliding window Stat. No $D$
Extreme learning machine (AnDi-ELM)[111] Sliding window ML Yes model, $\alpha$
Recurrent neural networksa [128] Sliding window ML Yes $\alpha$
Recurrent neural networksb (RANDI) [129] Sliding window ML Yes model, $\alpha$
Back-propagation neural network with MSD analysis[140] Sliding window ML Yes model
Extracting features of subtrajectories with historical experience learning (Deep-SEES)[141] Sliding window ML Yes model
Multi-feature random forest trajectory analysis[148] Sliding window ML Yes model
Bayesian model selection to hidden Markov modeling [46] Sliding window ML+Stat. No model, $v$
Gamma mixture and hidden Markov model[149] Sliding window ML+Stat. No $D$
Feature-engineered classical statistics combined with supervised deep learning (CONDOR)[150] Sliding window ML+Stat. Yes model, $\alpha$
Deep learning followed by moment scaling spectrum analysis (DL-MSS)[151] Sliding window ML+Stat. Yes model
Geometric deep learning framework with attention mechanisms (MAGIK)[135] Point-wise ML Yes model, $\alpha$, $D$
Deep convolutional networks (U-AnDi)[138] Point-wise ML Yes model, $\alpha$
Point-wise diffusion property prediction (STEP)[146] Point-wise ML Yes $D$, $\alpha$
Recurrent neural networksc[152] Point-wise ML Yes model
Nonparametric Bayesian inference (NOBIAS)[130] Point-wise ML+Stat. Yes model
[…]
00footnotetext: Note: Meanings of new symbols and abbreviations in the table are listed as follows. $v$: velocity. Stat.: statistics. Sim.: simulation. Seg.: segmentation.
To mitigate issues related to fixed segment lengths, researchers have developed the point-wise method, which is an approach inspired by the semantic segmentation in computer vision [153]. Semantic segmentation, also known as pixel-wise segmentation in image processing, classifies each pixel into specific categories. This technique enhances object recognition and delineates boundaries within images with high precision. Analogously, the point-wise method in trajectory analysis considers each data point along a trajectory as a potential changepoint. This allows for a more fine-grained and adaptable identification of state transitions, offering a distinct advantage in scenarios where diffusion states may change rapidly or irregularly.
One representative application of the point-wise technique is the STEP method developed by B. Requena et al.[146]. As shown in figure 6(a), this method leverages deep learning models, including CNN and Transformer[81] encoder, to segment the trajectory in a point-wise manner. This design allows the model to extract critical diffusion features such as diffusion coefficients and diffusion exponents at each time point. The STEP model is trained using the simulation-based method and validated with experimental data. This includes segmenting the trajectories of the integrin $\alpha 5\beta 1$ and pathogen recognition receptor DC-SIGN, both exhibiting varying diffusion coefficients. Notably, STEP demonstrates excellent performance in segmenting these complex trajectories, effectively identifying states with different diffusion coefficients [figure 6(b)].
Another example of point-wise method is the U-AnDi model proposed by X. Qu et al. [138]. As illustrated in figure 6(c), this model combines dilated causal convolutions to capture long-range dependencies, gated activation units for time-series information filtering, and a U-Net architecture to achieve fine-grained segmentation performance. Similar to the STEP method, U-AnDi predicts diffusion parameters for each point along a trajectory, including the diffusion exponent and diffusion model. As shown in figure 6(d), after training on simulated datasets, U-AnDi is further tested on experimental data involving the diffusion of transmembrane proteins on cell membrane surfaces. These experimental trajectories include two diffusion models: CTRW and FBM, with U-AnDi demonstrating excellent segmentation performance for both. In addition to the STEP and U-AnDi models, other machine learning approaches like RNN [152] and GNN [135] also employ point-wise methods to segment trajectories effectively. Detailed information regarding these models is listed in Table 1.
2.2.3 Further discussions
Both the sliding window and the point-wise method are effective in segmenting heterogeneous diffusion dynamics, yet each has its limitations. For the sliding window method, the selection of window length crucially impacts the model’s recall in detecting short-term changes. In contrast, the point-wise method benefits from its per-point prediction characteristic, which allows for more effective captures of transient states. However, every point along a trajectory can potentially be identified as a changepoint when employing the point-wise method. This can result in an increased incidence of false positives, reducing the precision of changepoint detection under certain conditions. Therefore, effective post-processing of predictions to filter out false changepoints is crucial for enhancing the segmentation performance of the point-wise method. Potential strategies to mitigate these false positives include the implementation of statistical tests to confirm the significance of potential changepoints, or the application of smoothing algorithms that help discern genuine changepoints from noise. Additionally, extra machine learning models can be trained to learn the typical patterns of changepoints in specific contexts, thereby reducing the likelihood of false detections. Although this post-processing issue has been partially addressed in the studies mentioned [146, 138], it remains a topic that requires continued research.
2.3 Anomalous Diffusion Challenge
Considering the proliferation of machine learning models aimed at analyzing anomalous diffusion, a standardized framework for evaluating their effectiveness remains essential. In the field of machine learning, benchmarking models against established datasets, like ImageNet [154] for image classification or COCO [155] for object detection, is the standard practice. Similarly, for the analysis of anomalous diffusion, establishing robust benchmarks is crucial to objectively assess and compare the performance of different models. This has led to the adoption of competitive platforms, similar to Kaggle [156], where machine learning models contend under uniform conditions using predefined tasks and datasets. In this section, we introduce the Anomalous Diffusion (AnDi) Challenge [89], which is a machine learning competition specifically designed for the analysis of anomalous diffusion trajectories.
The AnDi Challenge is organized by G. Muñoz-Gil et al. and is hosted on the CodaLab platform [157]. This competition features two distinct versions. The first iteration in the year 2020 focuses primarily on the inference of diffusion parameters, including the diffusion exponent and diffusion model [29]. While this version includes a segmentation task, its simplistic design does not fully replicate the complexities encountered in real-world scenarios. Consequently, the subsequent iteration of the challenge in the year 2024 shifts its emphasis towards the segmentation of heterogeneous diffusion dynamics [158], aiming to address this complex problem more directly and effectively. Details of the AnDi Challenge are listed in Table 2, including the task descriptions, evaluation metrics, number of teams for each task, and links to related works. The ranking information can be found on the official competition website [89] and is not listed in the table here for brevity. In the following, we provide an overview of each version of the AnDi Challenge, focusing on their specific objectives and respective datasets:
Figure 7: (a) Schematic diagram of three tasks in the $1^{\rm st}$ AnDi Challenge, from [29]; (b) Illustration of the two competition tracks and corresponding two subtasks in the $2^{\rm nd}$ AnDi Challenge, from [158].
Table 2: Detailed information of the AnDi Challenge.
Version Task description Metric Team number Related works
$1^{\rm st}$ Task 1: Inference of the diffusion exponent. Participants are required to infer the diffusion exponent of trajectories. Task 2: Classification of the diffusion model. Participants need to classify the trajectories into five diffusion models: ATTM, CTRW, FBM, LW, and SBM. Task 3: Two-state trajectory segmentation. The goal is to identify the single changepoint in each trajectory where the diffusion state changes just once, and then determine the diffusion model and exponent for each segment. Task 1: MAE Task 2: F1-Score Task 3: MAE, F1-Score, RMSE Task 1: 13 teams Task 2: 14 teams Task 3: 4 teams [29, 111, 115, 118, 123, 124, 125, 129, 131, 133, 150, 159, 160, 161, 162, 163, 164]
$2^{\rm nd}$ This challenge is organized along two tracks: Video track: For analysis of raw SPT videos. Trajectory track: For the analysis of trajectory data. Each track has two tasks: Task 1 (Single-trajectory task): For each trajectory, participants are supposed to list the changepoints. After that, they need to determine the diffusion coefficient, diffusion exponent, and motion type for each segment. Task 2 (Ensemble task): For each experimental condition, participants are required to provide the model type, number of states, mean and standard deviation of diffusion coefficients, and diffusion exponents for each state. Task 1: JSC, RMSE, MSLE, MAE, F1-Score Task 2: Wasserstein distance Video track: Task 1: 3 teams Task 2: 3 teams Trajectory track: Task 1: 18 teams Task 2: 11 teams [158]
00footnotetext: Note: New abbreviations in the table are listed as follows: RMSE: Root mean squared error; JSC: Jaccard similarity coefficient; MSLE: Mean squared logarithmic error.
  1. 1.
    As shown in figure 7(a), the $1^{\rm st}$ AnDi Challenge contains three tasks: the inference of diffusion exponent (Task 1), classification of diffusion model (Task 2), and segmentation of two-state trajectories (Task 3). Each task is further divided by the space dimensionality (1D, 2D, and 3D), resulting in nine independent subtasks. The trajectory data is generated by the Python package andi_datasets [165], based on five mathematical diffusion models: Annealed transient time motion (ATTM) [166], CTRW [41, 42], FBM [43], LW [56], and Scaled Brownian motion (SBM) [167]. By setting specific parameters in diffusion models, this Python package can generate synthetic diffusion trajectories with known labels. Participants can train their machine learning models using these simulated trajectories. A shared test set is then used by organizers to objectively evaluate the performance of different models, where the results are summarized in Ref. [29].
  2. 2.
    The $2^{\rm st}$ AnDi Challenge is organized along two tracks: video track and trajectory track. The video track focuses on processing raw SPT videos, while the trajectory track deals with the segmentation of trajectory data. As presented in figure 7(b), each track has two tasks: the single-trajectory task (Task 1), which involves segmenting individual trajectories, and the ensemble task (Task 2), which aims to estimate global parameters used in trajectory simulations. Here, we specifically detail the trajectory track, which is directly relevant to this article. To accurately replicate real experimental conditions of heterogeneous diffusion dynamics, the trajectory track’s dataset and objectives have been structured to reflect a higher level of complexity. Thus, to simplify the competition tasks, participants are required to solely process 2D trajectory data. These 2D trajectories are simulated based on five experimental models, namely single-state, multi-state, dimerization, transient-confinement, and quenched-trap models, which can emulate real anomalous diffusion experimental systems. By selecting experimental models and setting parameters, andi_datasets can generate a large number of simulated trajectories and corresponding labels. Utilizing these data, participants can train their models to segment trajectories and infer experimental parameters. The final results are summarized in Ref. [158] and will be officially published soon.
As subsequent versions continue to be refined, the AnDi Challenge can potentially establish crucial benchmarking protocols for the objective assessment of methods regarding the analysis of anomalous diffusion. Compared to the ImageNet dataset for image classification models, the AnDi Challenge aims to provide a similarly impactful framework for evaluating related models. This challenge facilitates the comparison of diverse methods under uniform testing conditions, which is vital for advancing the field towards robust and generalized solutions.
However, despite its significant contributions, the AnDi Challenge highlights a critical gap in this field: the reluctance within the scientific community to share real experimental data openly. Numerous studies on the machine learning characterization of anomalous diffusion rely heavily on the simulation-based method due to the lack of accessible real datasets. This situation constrains the development of machine learning tools that could be effectively applied to actual experimental analysis. To address this issue, there is a pressing need for the SPT research community to actively promote data sharing. Encouraging the open source of real experimental data would not only validate the applicability of these models in practical settings, but also accelerate the discovery of novel insights into the dynamics of anomalous diffusion processes.
3 Representation Learning of Anomalous Diffusion
Moving beyond the application-focused tasks of characterizing anomalous diffusion trajectories, a deeper understanding of representation learning, i.e., how to effectively represent the anomalous diffusion, emerges as a critical foundational step in the analysis of such complex phenomena. Effective representation of anomalous diffusion is crucial not only for enabling practical tasks like classification, segmentation, and parameter estimation, but also for embedding the intricate diffusion dynamics into a comprehensible and operational framework. The diversity and complexity of physical mechanisms underlying different types of anomalous diffusion necessitate versatile representation strategies. A robust representation of trajectory data, therefore, should distill the essence of corresponding diffusion dynamics into a form that is informative, discriminative, and adaptable to various analytical challenges.
To address this issue, we introduce three distinct strategies for representing the characteristics of anomalous diffusion in this section. The first involves the combination of predefined and handcrafted features to encapsulate the essential physical information of diffusion processes. The second employs neural networks to derive feature vectors from the penultimate layer, capturing a data-driven representation of diffusion dynamics. Lastly, we discuss the strategy of using latent representations through autoencoder (encoder-decoder) architectures [78], which not only reflect the underlying mechanisms but also facilitate the reconstruction of trajectories.
3.1 Diffusion fingerprint: Combination of predefined features
Prior to the widespread integration of machine learning into the study of anomalous diffusion, researchers have already employed a range of physical parameters to characterize the diffusion dynamics, such as MSD and diffusion exponent. In contrast to machine learning methods that autonomously extract features from data, the derivation of these parametric features necessitates a deep understanding of the physical principles underlying anomalous diffusion. Consequently, they are typically predefined and manually extracted, anchored firmly in domain-specific knowledge. Therefore, the combination of these predefined features not only captures the complex behaviors inherent in anomalous diffusion but also provides substantial interpretability.
Table 3: Overview of predefined feature combinations used in representative studies to characterize anomalous diffusion through feature-based models. In Ref. [112], the authors evaluate the performance of three different feature combinations separately, designated as Set A, Set B, and Set C.
Random forest [112] LR [168] XGBoost [110]
Set A Set B Set C
Anomalous exponent
Diffusion coefficient
MSD ratio
Efficiency
Straightness
Empirical VACF
Maximal excursion
p-variation-based statistics
Asymmetry
Fractal dimension
Gaussianity
Kurtosis
Trappedness
Standardised maximum distance
Exponent of power function
Residence times in each state
Average residence time
Average step length
Quality of power law fit
Intermediate time spread
Number of points in a trajectory
Mean maximal excursion
D’Agostino-Pearson test statistic
Kolmogorov-Smirnov statistic
Noah exponent
Moses exponent
Joseph exponent
Detrending moving average
Average moving window characteristics
Maximum standard deviation
[…]
00footnotetext: Note: Here, LR is the abbreviation of logistic regression.
As discussed in section 2.1.1, feature-based machine learning models successfully leverage these predefined features to infer diffusion parameters. Nevertheless, selecting the most effective combination of predefined features to represent anomalous diffusion accurately remains a central challenge. Extensive research employing feature-based models has explored a diverse array of feature combinations to effectively characterize the diffusion dynamics. For an intuitive presentation, we summarize the feature combinations used in three representative works [112, 168, 110] in Table 3. As we see in this table, the selected features encompass not only specific anomalous diffusion parameters such as the diffusion exponent and MSD but also standard statistical measures like kurtosis and standard deviation. This composition of predefined features is strategically crafted to represent diffusion trajectories within an established knowledge framework. It maximizes the informational content while ensuring that the features maintain clear physical meanings and robust interpretability.
Figure 8: Schematic diagram of the diffusion fingerprint for representing and analyzing the single-particle tracking trajectory data, from [168].
Building upon this framework, as depicted in figure 8, H. D. Pinholt et al.[168] advance the field of representation learning by introducing the term “diffusion fingerprint" to describe the unique characteristics of combined features. This innovative concept underscores the precision with which these feature combinations can delineate the dynamic profiles of different diffusion types, serving a role analogous to how a fingerprint distinctively identifies an individual. However, as highlighted in section 2.1.1, the study by H. Loch-Olszewska et al.[112] emphasizes that no single feature combination excels across all types of anomalous diffusion. This observation suggests that there are inherent limitations in the representational power of predefined feature combinations when characterizing the properties of anomalous diffusion.
Therefore, to enhance the representational capacity of feature combinations while maintaining their high interpretability, further exploration of new and effective diffusion or statistical features beyond those already known is necessary. For example, M. Mangalam et al.[169] have advanced the analysis of non-ergodic diffusion processes by incorporating fractal and multifractal descriptors that capture ergodic properties. By employing the Chhabra-Jensen method [170], they calculate the multifractal spectrum curve, which relates the singularity exponents of these descriptors to the Hausdorff dimension. This approach effectively illustrates the nonlinearity and complexity inherent in different types of diffusion. For instance, the multifractal spectrum of FBM typically appears symmetric, reflecting relatively consistent fluctuations throughout the process. Conversely, SBM exhibits an asymmetric singularity spectrum, indicating fluctuations concentrated at particular scales. Leveraging the discriminative capabilities of the multifractal spectrum, H. Seckler et al.[171] further select nine spectrum-related parameters to form a multifractal spectral (MFS) feature set, which serves as a novel representation for diffusion processes. By integrating these MFS features with the predefined features listed in the XGBoost column of Table 3 , the classification performance of machine learning models is substantially enhanced. This integration underscores the effectiveness of MFS features in enriching the diffusion feature landscape, highlighting the importance of mining new features for the analysis of anomalous diffusion.
3.2 Data-driven representation learning by neural networks
The existing frameworks of knowledge do not completely capture all the nuanced features of diffusion processes, inherently imposing an upper limit on feature discovery based on traditional understandings. This limitation becomes glaringly apparent in efforts to characterize complex diffusion trajectories. However, inspired by data-driven machine learning approaches, enabling models to autonomously learn from data to extract effective representations could transcend these traditional constraints. This shift would free the analysis from the constraints of existing theoretical frameworks, allowing for the capture of high-level abstract representations of diffusion dynamics that are more effective and encompassing.
This data-driven representation learning, extensively applied in fields such as image, text, and speech analysis, involves transforming raw data into embedded vectors via deep neural networks[172, 173]. By constructing databases of these embeddings, the approach significantly enhances the efficiency of data retrieval, matching, and classification tasks. Therefore, the representation of anomalous diffusion can also benefit from this data-driven approach, wherein deep neural networks autonomously perform representation learning directly from trajectory data. Here, we introduce two commonly utilized data-driven representations of diffusion processes: feature vector from the penultimate layer of neural network, and latent representation from the autoencoder.
3.2.1 Feature vector from the penultimate layer of neural network
In general, when employing a neural network for parameter inference, the final prediction is determined by the linear operations in the last fully connected layer. In particular, these simple operations are only applied to the output from the penultimate layer, indicating that the model performance is substantially dependent on the representational power of this output. Similarly, in high-performing neural networks tailored for anomalous diffusion analysis, the feature vector from the penultimate layer encapsulates significant information that delineates the diffusion dynamics. This comprehensive content is vital for the model’s robust performance, making this feature vector a potent representation of the diffusion process.
Figure 9: (a) UMAP visualization of feature vectors of diffusion trajectories from WADNet across five different diffusion models and varying lengths, from [123]; (b) UMAP representation of trajectory feature vectors from the GNN model. Each plot, arranged from left to right and from top to bottom, focuses on the following aspects respectively: diffusion model, diffusion exponent, trajectory length, mixed trajectories, BM, and OU process, from [133].
However, representations derived from data-driven methods typically lack explicit physical or statistical meaning, rendering them less interpretable. Despite this limitation, it is still possible to gain insights into how these feature vectors represent anomalous diffusion characteristics by examining their distributions through visualization techniques. Dimensionality reduction techniques like t-distributed stochastic neighbor embedding (t-SNE) [174] and uniform manifold approximation and projection (UMAP) [175] allow for the visualization of high-dimensional feature vectors in a lower-dimensional space. For example, D. Li et al.[123] utilize UMAP to visualize feature vectors from WADNet on a 2D plane, as illustrated in figure 9(a). This visualization specifically represents feature vectors of trajectories across five different diffusion models and varying lengths. It is evident that as the trajectory length increases, trajectories from different diffusion models become more distinctly separated. This result demonstrates that the representational ability of feature vectors to differentiate among these diffusion models becomes more pronounced with increasing trajectory lengths. In addition, the cluster distribution distinctly separates the motion patterns of LW from the others, while FBM and SBM show overlapping behaviors, as do ATTM and CTRW. This distribution align with the theoretical definitions of these diffusion models, indicating that the feature vectors from the penultimate layer of WADNet effectively capture the characteristics of diffusion trajectories.
Another work that highlights the feature vectors from neural networks is by H. Verdier et al. [133], where these vectors are extracted by a GNN model. As illustrated in figure 9(b), the UMAP visualizations capture the diffusion dynamics across various parameters such as diffusion model, diffusion exponent, and trajectory length. The clustering clearly demonstrates the GNN’s ability to discern differences among these parameters, underscoring the effectiveness of feature vectors in capturing the essence of diffusion processes. Additionally, the authors assess the robustness and generalizability of these representations by testing mixed trajectories and unseen diffusion models. For mixed trajectories from FBM and ATTM models, the feature vectors exhibit a smooth transition across different types of motion as their proportions vary, effectively interpolating between the characteristics of these two diffusion models. Further, feature vectors from trajectories generated using diffusion models not included in the training set are incorporated, specifically Brownian motion (BM) and the Ornstein–Uhlenbeck (OU) process [176]. Corresponding distribution of clusters show that these trajectories overlap significantly with those of SBM and FBM. This consistency with theoretical expectations confirms the capacity of feature vectors from the penultimate layer of neural network to generalize and robustly represent diverse diffusion dynamics.
3.2.2 Latent representation from the autoencoder
Figure 10: (a) Schematic diagram of the autoencoder architecture, from [177]; (b) Distribution of latent neurons activations, showing clear relations to the diffusion parameters, from [177]; (c) Illustration of the experimental system of gold nanorods, where both translational (top) and rotational (bottom) trajectories are effectively denoised by the LSTM-VAE model, from [141].
While leveraging trained neural networks for parameter inference is a highly effective and efficient method for extracting representations of diffusion trajectories, it is important to note that these models are specifically designed for inference tasks. This specialized focus may inadvertently lead to the neglect of certain trajectory features that do not directly influence the target parameters. As a result, the derived representations can exhibit inherent biases, capturing a skewed perspective of diffusion dynamics. Therefore, developing neural networks that focus primarily on the extraction of representations is crucial for advancing the representation learning of anomalous diffusion.
For that purpose, researchers have shifted their emphasis towards machine learning models based on the autoencoder architecture. This architecture is particularly adept at distilling high-dimensional data into efficient latent representations. As show in figure 10(a), the autoencoder consists of two components: an encoder and a decoder. The encoder compresses the input data into a low-dimensional hidden layer, which forms the latent representation. The decoder then attempts to reconstruct the original input data from this latent space. It is evident that the training of autoencoder models does not rely on data labels, making them considerably more suitable for representation learning than discriminative models that require labeled data.
The pioneering work in this area is conducted by G. Muñoz-Gil et al.[178], where an autoencoder model is utilized for anomaly detection in an unsupervised way. Although this work does not specifically highlight the representation learning of anomalous diffusion, their application of autoencoder and realization of trajectory reconstruction has significantly inspired subsequent research in this domain. Following this work, G. Fernández-Fernández et al.[177] introduce a $\beta$-variational autoencoder ($\beta$-VAE) tailored for the representation learning of stochastic processes. Besides extracting latent representations and reconstructing trajectories, the authors also try to identify the minimal physical parameters that accurately describe specific stochastic behaviors. By constraining the number of latent neurons, they compel the $\beta$-VAE model to encode only the variables related to the physical process, such as diffusion coefficient and diffusion exponent, in the latent space. This constraint ensures that the remaining neurons exhibit a pronounced correlation with these key physical variables [figure 10(b)], thus substantially enhancing the interpretability of the representations and enabling the autonomous discovery of unknown parameters.
On the other hand, autoencoders excel in processing experimental trajectories directly due to their independence from labeled data. This attribute enables versatile applications in managing raw data, particularly effective in denoising experimental trajectories through latent representation and trajectory reconstruction. For example, the Deep-SEES method, developed by Y. Zhang et al.[141], utilizes the LSTM-VAE model to extract latent representations from experimental trajectories. This approach enables the reconstruction of these trajectories, effectively denoising and preserving their physical characteristics. The robustness of these reconstructions is tested across challenging experimental conditions, including trajectories of gold nanorods in active enzyme systems and protein liquid-liquid phase separation systems. These trajectories are characterized by high noise levels due to low photon flux during data acquisition. As highlighted in figure 10(c), the LSTM-VAE model effectively denoises these trajectories for both translational and rotational movements, maintaining their physical accuracy and facilitating the subsequent classification of different motion types. The results not only demonstrate the exceptional performance of autoencoders in denoising but also affirms the potent representational capability of latent representations in capturing the dynamics of diffusion processes.
3.3 Further discussions
Table 4: Comparison of three methods for the representation learning of anomalous diffusion.
Predefined feature Feature vector Latent
Base of representation Known knowledge Data Data
Representational power Moderate Strong Strong
Interpretability Strong Weak Weak
Dependence on labels Weak Strong Weak
Computational demand Low Moderate High
Automatic feature discovery Incapable Able Able
Ability to reconstruct Incapable Incapable Able
[…]
00footnotetext: Note: Here, “predefined feature" refers to the combination of predefined features, “feature vector" to the feature vector from the penultimate layer of neural network, and “latent" to the latent representation from the autoencoder.
As detailed in the preceding sections, each of the three methods for the representation learning of anomalous diffusion specializes in distinct aspects, highlighting their unique contributions to understanding diffusion processes. For applications where interpretability is paramount, such as in scientific research where understanding the mechanisms of diffusion is crucial, predefined features might be preferred. In contrast, for applications where model performance and adaptability to complex data are more critical, feature vectors from neural networks and latent representations from autoencoders may offer superior solutions. To provide a clearer overview, we have compared the three methods across multiple dimensions in Table 4, offering valuable insights into choosing the most appropriate representation technique for analyzing diffusion processes. Future research may consider the use of hybrid models that leverage both data-driven features and domain-specific knowledge. This strategy could harness the collective strengths of these two types of representations, potentially establishing enhanced benchmarks for the analysis of anomalous diffusion.
4 Conclusions
In summary, the integration of machine learning techniques in the analysis of anomalous diffusion is systematically introduced in this paper. We have discussed several key areas in this field, including the single trajectory characterization via machine learning and the representation learning of anomalous diffusion. For the characterization of single trajectory properties, machine learning techniques demonstrate superior performance compared to traditional methods in inferring diffusion parameters and segmenting trajectories. Moreover, the variety of machine learning approaches, spanning from classical machine learning to advanced deep learning, provides an extensive array of options that can be customized to meet the distinct demands of different scenarios. Through machine learning competitions such as the AnDi Challenge, the field has pushed towards standardized benchmarks that facilitate the evaluation and comparison of various analytical models under uniform conditions.
Regarding the representation learning of anomalous diffusion, we introduce and compare three principal strategies: the combination of predefined features, the feature vector from the penultimate layer of neural network, and the latent representation from the autoencoder. Each strategy has its merits and limitations, suggesting that the choice of method should be aligned with specific research needs. Predefined features, while limited in adaptability, offer significant interpretability, making them invaluable in contexts where understanding the underlying physical mechanisms is paramount. On the other hand, data-driven representations, including feature vectors and latent representations, provide powerful tools for handling more complex trajectories, although they often lack the actual physical meanings.
Despite substantial advances, this field still faces significant challenges, particularly in terms of data availability, encompassing both simulated and experimental datasets. Currently, the simulated data available generally do not fully replicate the complexities of real-world trajectories, with many diffusion models lacking corresponding simulation tools. This restriction often confines research to diffusion models included in available tools such as andi-datasets. Thus, increasing the availability of trajectory simulation software is essential, which would significantly enhance the capabilities of machine learning models trained via simulation-based methods. On the other hand, the access to experimental trajectory data is largely restricted, which hinders the involvement of machine learning experts in anomalous diffusion research. Therefore, the establishment of an open-source ecosystem for sharing experimental data is important, requiring collective efforts from the research community to accelerate development in this domain.
Moving forward, machine learning analysis of anomalous diffusion presents two promising research directions. The first involves the creation of innovative methods and models to address complex analytical challenges in trajectory analysis, such as accurate changepoint detection and forecasting of random walks. This direction necessitates a focus on advanced techniques for time-series analysis, potentially leveraging cutting-edge technologies like large language models [86]. The second emphasizes enhancing the interpretability of machine learning models to facilitate a deeper understanding of the physical underpinnings of diffusion processes. This will require embedding domain-specific knowledge of anomalous diffusion into machine learning architectures to enable more insightful analysis and interpretation. In conclusion, leveraging machine learning to investigate anomalous diffusion represents a vital approach that is reshaping our understanding of complex systems. This progressive integration not only promises to refine our predictive capabilities but also holds potential to broaden our theoretical frameworks of anomalous diffusion, thereby offering profound insights into the complex behaviors and interactions that drive diffusion phenomena across physical, biological, and technological domains.
[…]
Acknowledgements
We thank Yu Zhang for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12104147) and the Fundamental Research Funds for the Central Universities.